OpenStudy (yoyo2443):

Solve the following system of linear equation using any algebraic method. If possible check your solution.

OpenStudy (yoyo2443):

\[6y=15-9x\] \[6x=10-4y\]

OpenStudy (dumbcow):

First rewrite in standard form (Ax + By = C) 9x + 6y = 15 6x + 4y = 10 Next factor out greatest common factor from each equation 3x + 2y = 5 3x +2y = 5 Notice they are now the exact same equation This means they are the same line so the 2 lines intersect at all points. Infinite solutions

OpenStudy (yoyo2443):

Oooh alright, so there isn't any exact solution for it?

OpenStudy (dumbcow):

No every possible point is a solution

OpenStudy (yoyo2443):

oh right right, thanks!